OpenStudy (anonymous):

Viola drives 4 kilometers up a hill that makes an angle of 8 degrees with the horizontal. To the nearest tenth of a kilometer, what horizontal distance has she covered?

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (solomonzelman):



OpenStudy (anonymous):

So i use Cos?

OpenStudy (anonymous):

or 4000*cos (8)

OpenStudy (solomonzelman):

Have you learned the law of sines yet ? A I . I . I . b c I . I . I . I___________. B a C \[\frac{sinA}{a}=\frac{sinB}{b}=\frac{sinC}{c}\]

OpenStudy (anonymous):

no i didnt that isnt even what the lesson is on

OpenStudy (solomonzelman):

OKAY... what is your lesson on ? What DID you learn in the lesson ?

OpenStudy (anonymous):

I dont know but im getting fustrated